# DiffTester: Accelerating Unit Test Generation for Diffusion LLMs via Repetitive Pattern

### v1

@@ 1 Introduction
Unit testing plays a vital role in software development, ensuring that a functionally discrete program unit (e.g.,a method) behaves correctly and meets the intended design expectations Olan (2003); Gren & Antinyan (2017). However, manually writing high-quality Unit Tests (UTs) is often extremely time-consuming and labor-intensive Runeson (2006), especially in large-scale software development scenarios Shang et al. (2025). To reduce the manual burden, recent research has increasingly sought many automated approaches to generate UTs Chipounov et al. (2011); Tufano et al. (2020); Fraser & Arcuri (2011b). Among these, advanced Large Language Models (LLMs) have rapidly become the mainstream solution for automated Unit Test Generation (UTG), due to their strong code understanding and code generation abilities Wang et al. (2024); Bhatia et al. (2024).
@@
In large-scale software development projects, it is often necessary to generate a substantial number of UTs, sometimes reaching hundreds or thousands Robinson et al. (2011); Li et al. (2006). However, most existing LLMs Hui et al. (2024); Achiam et al. (2023) generate UTs one token at a time during each forward pass, which substantially increases both time consumption and computational cost, leading to inefficient UTG Yang et al. (2024). Fortunately, emerging Diffusion Large Language Models (dLLMs) Nie et al. (2025b); Ye et al. (2025) exhibit strong potential in UTG and other code-related tasks Li et al. (2025a); Khanna et al. (2025), which adopt a unique generation paradigm that naturally enables multi-token prediction and flexible generation order. Specifically, at each inference step, dLLMs predict a candidate token for every [MASK] position and then remask those with relatively low confidence according to a remasking strategy.
@@
Although dLLMs are theoretically capable of generating multiple tokens in each forward pass, existing studies Zeng et al. (2025); Barr et al. (2014) typically set the number of tokens generated per step to a very small value (e.g.,one or two). As a result, their inference speed is often comparable to that of non-diffusion LLMs, which fail to fully exploit the efficiency advantages of dLLMs in theory Li et al. (2025a). Our preliminary experiments further reveal that when the number of tokens generated per step is increased, the quality of the generated UTs degrades sharply, and in many cases, even the syntactic correctness of the test cases cannot be guaranteed (for more details, see Appendix C.2.1). This limitation greatly hinders the efficiency of dLLMs in UTG, highlighting the need for an acceleration framework specifically tailored to dLLMs for UTG.
@@
To accelerate dLLMs in UTG, we conduct an in-depth analysis of the characteristics of this task. We find that for the same focal method, the generated UTs often share repetitive patterns, frequently exhibiting substantial structural and syntactic repetition. For instance, two generated unit tests often share similar patterns and differ only in specific details. From the perspective of their Abstract Syntax Trees (ASTs) Peacock et al. (2021), these differences typically appear only in certain leaf nodes or low-level non-leaf nodes. Based on this insight, we propose DiffTester to accelerate unit test generation for diffusion LLMs via repetitive pattern. Specifically, we first prompt dLLMs within a single batch to produce multiple UTs for a given focal method. At selected steps of the generation process, we parse the generated code into multiple ASTs and extract the common nodes across them, which we regard as indicative of patterns inherent to the test cases for the focal method. The tokens corresponding to these common nodes are then generated in a single step. In this way, DiffTester dynamically and appropriately increases the number of tokens generated at each step, while preserving the quality of the resulting test cases.
@@
We perform extensive experiments to validate the effectiveness of our proposed DiffTester on the TestEval benchmark Wang et al. (2025) (whose focal methods are all implemented in Python) with two representative dLLMs, including Dream Ye et al. (2025) and DiffuCoder Gong et al. (2025). To enable comprehensive evaluation and to avoid potential data contamination Deng et al. (2023), we further extend the TestEval benchmark to multiple programming languages, resulting in TestEval-C++ and TestEval-Java. Experimental results demonstrate that, under the same runtime or computational budget, applying DiffTester substantially improves coverage. From another perspective, to achieve the same level of coverage, DiffTester can often reduce both time and computational cost by more than half. We believe that DiffTester effectively accelerates dLLMs in UTG and can contribute to software quality assurance.
@@ 2.1 Inference Process of dLLMs
In this section, we introduce how mainstream dLLMs, such as Dream Ye et al. (2025) and LLaDa Nie et al. (2025b), perform inference. These models progressively refine a sequence consisting of $L$ special [MASK] tokens into the final generated output (e.g., some test cases). Formally, let $\mathcal{Y}^{0}=(y_{i}^{0})_{i=1}^{L}$, where $y_{i}^{0}=\texttt{[MASK]}$, denote the initial fully masked sequence. Given a prompt $p$, the dLLM performs $T$ steps, each consisting of a forward pass, and eventually produces the output $\mathcal{Y}^{T}=(y_{i}^{T})_{i=1}^{L}$, which contains no remaining [MASK] tokens.
At each step $t$, the dLLM performs one forward pass to predict the probability distribution for every position that is filled with a [MASK] token in parallel. For each such position $i$, the model then samples a non-[MASK] token $\hat{y}_{i}^{t}$ from its corresponding probability distribution, a process analogous to the sampling procedure used in autoregressive LLMs. After this step, the model proceeds with the remasking operation Nie et al. (2025b); Ye et al. (2025), where a confidence score is assigned to each predicted token (e.g., based on entropy, maximum probability of the distribution, or by randomly specifying a value). Tokens with the highest confidence are retained, for instance, by selecting the top-$k$ predictions or by keeping those whose confidence exceeds a predefined threshold Wu et al. (2025b); Wei et al. (2025b), while other tokens are reverted back to [MASK] for further refinement in subsequent steps. Many prior studies Li et al. (2025a); Wu et al. (2025a) have shown that increasing the number of tokens generated at each step can significantly degrade the quality of the outputs.
@@ 2.2 Dive into Unit Tests Generation
Given a specific focal method, unit test generation typically needs to produce multiple test cases, each consisting of a corresponding test prefix and a test assertion Lemieux et al. (2023); Zamprogno et al. (2022). The test prefix is primarily used to construct the testing data and environment, whereas the test assertion serves to verify the correctness of the focal method Yuan et al. (2024).
@@ 3.1 Overview of DiffTester
At one step $t$, for the $k$-th instance in the batch, the model first predicts tokens for all [MASK] positions, producing an intermediate response $\mathcal{Y}^{t,k}=(\hat{y}_{i}^{t,k})_{i=1}^{L}$. We then follow the standard remasking strategy defined in existing dLLMs (e.g., selecting the top-$k$ tokens according to their confidence scores) to determine which tokens will be retained Ye et al. (2025); Xie et al. (2025); Gong et al. (2025). Subsequently and most importantly, we extract shared patterns across the generated unit test cases and leverage them to further retain additional tokens, with the detailed procedure presented in Section 3.2. After completing the procedure described above, we retain all selected tokens, while the remaining tokens at [MASK] positions are reverted back to [MASK], and the process proceeds to the next step.
@@ 3.2 Exploiting Repetitive Patterns to Accelerate UTG
We begin by analyzing the key challenges involved in leveraging such repetitive patterns, which lie in two main aspects. ❶ The first challenge is how to effectively extract shared patterns across multiple test cases. As illustrated in Figure 1, repetitive structural and syntactic patterns can be easily observed among different test cases for the same focal method. However, transforming these intuitive and abstract notions of repetitive patterns into a concrete representation-that is, determining which tokens are structurally repeated—remains a non-trivial problem. ❷ The second challenge is how to ensure sufficient diversity in the generated test cases. Achieving high test coverage relies heavily on diversity Peacock et al. (2021); Yang (2023), yet the process of extracting shared patterns may risk reducing it. For instance, the model might generate identical input data across multiple test cases. Designing a method that can exploit shared patterns while preserving the diversity of test data is therefore another important challenge.
@@
Extract Repetitive Patterns. We noticed Abstract Syntax Trees (ASTs) provide an objective representation of the syntactic structure of code Sun et al. (2023); Suttichaya et al. (2022), and each node in an AST can be accurately mapped to a specific set of tokens. Based on this property, our approach compares the ASTs of different test cases’ code and attempts to merge them as much as possible, which means identifying the nodes that are shared across different ASTs. When two or more ASTs can be merged into a non-empty tree, this indicates the existence of a repetitive pattern. We then locate the tokens in the intermediate test cases that correspond to the merged AST.
@@
Ensure Sufficient Diversity. The diversity of unit test cases largely depends on the variability of the input data constructed for the focal method, which in turn is primarily determined by literal values such as integers or floats Yang (2023). To preserve this variability, we exclude the AST nodes corresponding to such literal values from the merging process. This design ensures that even if two test cases construct similar data at an intermediate step, the corresponding tokens are not retained in a single step but are instead remasked for subsequent refinement, allowing higher diversity.
@@ 4.1 Experiment setup
Models. We evaluate our approach using two representative dLLMs, namely DiffuCoder-7B-cpGRPO Gong et al. (2025) and Dream-v0-Instruct-7B Ye et al. (2025). For brevity, we refer to them as DiffuCoder and Dream in the remainder of this paper.
Benchmarks. To evaluate the performance of DiffTester, we conduct experiments on the TestEval benchmark Wang et al. (2025). TestEval is specifically designed to assess the capability of models in unit test generation, and it comprises 210 Python programs collected from LeetCode. We denote this benchmark as TestEval-Python. To further validate the generalization ability of DiffTester across different programming languages and to avoid potential data contamination, we extend the TestEval-Python benchmark by incorporating two additional programming languages: C++ and Java. Following the construction methodology of TestEval-Python, we systematically gather corresponding C++ and Java implementations of the same set of 210 programs. We denote these benchmarks as TestEval-C++ and TestEval-Java, respectively.
@@ 4.2 Main Results
DiffTester does not compromise the maximum achievable test coverage. The highest test coverage attainable without time constraints is another important optimization target, particularly for scenarios such as vehicle control systems Conrad & Fey (2017); Zhang et al. (2024). We find that DiffTester has little to no negative impact on the maximum test coverage and in some cases even improves it. For example, as shown in Figure 4, on the TestEval-Python benchmark with the Dream model, applying DiffTester leads to a slight increase in max test coverage.
@@ 5.1 Comparison with dLLM Acceleration for General Task
We observe that several contemporary works have also explored training-free approaches to accelerate dLLMs Wei et al. (2025b); Ben-Hamu et al. (2025b); Israel et al. (2025). However, these approaches typically focus on general-purpose tasks, whereas DiffTester is designed specifically for unit test generation (UTG), a domain where efficiency is particularly critical. To highlight the difference, we compare DiffTester with a representative approach, EB-Sampler Ben-Hamu et al. (2025b). EB-Sampler accelerates dLLM inference through an Entropy-Bounded unmasking procedure, which dynamically unmasks multiple tokens under a predefined approximate error tolerance.
@@ 6.1 Unit Test Generation
Unit tests Olan (2003); Runeson (2006) are essential for verifying the behavior of program units and serve as a foundational mechanism for early fault detection and prevention in software development. However, writing unit tests manually is highly labor-intensive, which has motivated extensive research on automated unit test generation Dakhel et al. (2024); Aniche (2022). Traditional techniques primarily rely on search-based Delgado-Pérez et al. (2022); McMinn (2004), random-based Pacheco et al. (2007), constraint-based Csallner et al. (2008), and symbolic execution-based approaches Baldoni et al. (2016), with widely used tools such as EvoSuite Fraser & Arcuri (2011a), Pynguin Lukasczyk & Fraser (2022), and KLEE Cadar et al. (2008). Nevertheless, these traditional approaches often suffer from limitations such as low coverage Barr et al. (2014).
With the rise of large language models (LLMs) Achiam et al. (2023), recent advances have increasingly explored their application to automated test case generation. Representative approaches include ChatTester Yuan et al. (2024), TestPilot Schäfer et al. (2023), and ChatUnitTest Chen et al. (2024). Although LLMs have significantly improved test coverage, existing LLMs used for unit test generation almost exclusively adopt a token-by-token generation paradigm Li et al. (2025a), which constrains the efficiency of the unit test generation.
@@ 6.2 Diffusion Large Language Model
Diffusion language models have recently become a focal point in AI research Sahoo et al. (2024); Liang et al. (2025) and can be categorized into continuous and discrete formulations Li et al. (2025d). Empirical evidence Gong et al. (2024) suggests that discrete diffusion language models scale more effectively to larger model sizes, which has led to the development of diffusion large language models such as LLaDA Nie et al. (2025b) and Gemini Diffusion DeapMind (2025). Within this line of work, MMaDA Yang et al. (2025) and LLaDA-V You et al. (2025) extend dLLMs to the multimodal domain; TreeDiff Zeng et al. (2025) explores integrating dLLMs with the structural characteristics of code; DLLM-Cache Liu et al. (2025c) and Sparse-dLLM Song et al. (2025a) propose inference acceleration strategies based on KV-cache mechanisms; DiffuCoder Gong et al. (2025) and LLaDA-1.5 Zhu et al. (2025b) investigate reinforcement learning tailored to dLLMs.
@@ C.1.1 Models
We primarily conducted experiments using the dLLMs DiffuCoder-7B-cpGRPO (Gong et al., 2025) (DiffuCoder in this paper for short) and Dream-v0-Instruct-7B Ye et al. (2025) (Dream in this paper for short).

### v2

@@ 1 Introduction
Unit testing plays a vital role in software development, ensuring that a functionally discrete program unit (e.g.,a method) behaves correctly and meets the intended design expectations UT1; gren2017relation. However, manually writing high-quality Unit Tests (UTs) is often extremely time-consuming and labor-intensive UT2, especially in large-scale software development scenarios shang2025large. To reduce the manual burden, recent research has increasingly sought many automated approaches to generate UTs chipounov2011s2e; tufano2020unit; EvoSuite. Among these, advanced Large Language Models (LLMs) have rapidly become the mainstream solution for automated Unit Test Generation (UTG), due to their strong code understanding and code generation abilities wang2024software; bhatia2024unit.
@@
In large-scale software development projects, it is often necessary to generate a substantial number of UTs, sometimes reaching hundreds or thousands robinson2011scaling; li2006large. However, most existing LLMs hui2024qwen2; achiam2023gpt generate UTs one token at a time during each forward pass, which substantially increases both time consumption and computational cost, leading to inefficient UTG yang2024evaluation. Fortunately, emerging Diffusion Large Language Models (dLLMs) llada; dream exhibit strong potential in UTG and other code-related tasks zhang2025beyond; mercury, which adopt a unique generation paradigm that naturally enables multi-token prediction and flexible generation order. Specifically, at each inference step, dLLMs predict a candidate token for every [MASK] position and then remask those with relatively low confidence according to a remasking strategy.
@@
Although dLLMs are theoretically capable of generating multiple tokens in each forward pass, existing studies zeng2025treediff; time-consuming1 typically set the number of tokens generated per step to a very small value (e.g.,one or two). As a result, their inference speed is often comparable to that of non-diffusion LLMs, which fail to fully exploit the efficiency advantages of dLLMs in theory zhang2025beyond. Our preliminary experiments further reveal that when the number of tokens generated per step is increased, the quality of the generated UTs degrades sharply, and in many cases, even the syntactic correctness of the test cases cannot be guaranteed (for more details, see Appendix C.2.1). This limitation greatly hinders the efficiency of dLLMs in UTG, highlighting the need for an acceleration framework specifically tailored to dLLMs for UTG.
@@
To accelerate dLLMs in UTG, we conduct an in-depth analysis of the characteristics of this task. We find that for the same focal method, the generated UTs often share repetitive patterns, frequently exhibiting substantial structural and syntactic repetition. For instance, two generated unit tests often share similar patterns and differ only in specific details. From the perspective of their Abstract Syntax Trees (ASTs) peacock2021automatic, these differences typically appear only in certain leaf nodes or low-level non-leaf nodes. Based on this insight, we propose DiffTester to accelerate unit test generation for diffusion LLMs via repetitive pattern. Specifically, we first prompt dLLMs within a single batch to produce multiple UTs for a given focal method. At selected steps of the generation process, we parse the generated code into multiple ASTs and extract the common nodes across them, which we regard as indicative of patterns inherent to the test cases for the focal method. The tokens corresponding to these common nodes are then generated in a single step. In this way, DiffTester dynamically and appropriately increases the number of tokens generated at each step, while preserving the quality of the resulting test cases.
@@
We perform extensive experiments to validate the effectiveness of our proposed DiffTester on the TestEval benchmark testeval (whose focal methods are all implemented in Python) with two representative dLLMs, including Dream dream and DiffuCoder DiffuCoder. To enable comprehensive evaluation and to avoid potential data contamination deng2023investigating, we further extend the TestEval benchmark to multiple programming languages, resulting in TestEval-C++ and TestEval-Java. Experimental results demonstrate that, under the same runtime or computational budget, applying DiffTester substantially improves coverage. From another perspective, to achieve the same level of coverage, DiffTester can often reduce both time and computational cost by more than half. We believe that DiffTester effectively accelerates dLLMs in UTG and can contribute to software quality assurance.
@@ 2.1 Inference Process of dLLMs
In this section, we introduce how mainstream dLLMs, such as Dream dream and LLaDa llada, perform inference. These models progressively refine a sequence consisting of $L$ special [MASK] tokens into the final generated output (e.g., some test cases). Formally, let $\mathcal{Y}^{0}=(y_{i}^{0})_{i=1}^{L}$, where $y_{i}^{0}=\texttt{[MASK]}$, denote the initial fully masked sequence. Given a prompt $p$, the dLLM performs $T$ steps, each consisting of a forward pass, and eventually produces the output $\mathcal{Y}^{T}=(y_{i}^{T})_{i=1}^{L}$, which contains no remaining [MASK] tokens.
At each step $t$, the dLLM performs one forward pass to predict the probability distribution for every position that is filled with a [MASK] token in parallel. For each such position $i$, the model then samples a non-[MASK] token $\hat{y}_{i}^{t}$ from its corresponding probability distribution, a process analogous to the sampling procedure used in autoregressive LLMs. After this step, the model proceeds with the remasking operation llada; dream, where a confidence score is assigned to each predicted token (e.g., based on entropy, maximum probability of the distribution, or by randomly specifying a value). Tokens with the highest confidence are retained, for instance, by selecting the top-$k$ predictions or by keeping those whose confidence exceeds a predefined threshold wu2025fast; wei2025accelerating, while other tokens are reverted back to [MASK] for further refinement in subsequent steps. Many prior studies zhang2025beyond; wu2025fast have shown that increasing the number of tokens generated at each step can significantly degrade the quality of the outputs.
@@ 2.2 Dive into Unit Tests Generation
Given a specific focal method, unit test generation typically needs to produce multiple test cases, each consisting of a corresponding test prefix and a test assertion lemieux2023codamosa; zamprogno2022dynamic. The test prefix is primarily used to construct the testing data and environment, whereas the test assertion serves to verify the correctness of the focal method ChatTester.
@@ 3.1 Overview of DiffTester
At one step $t$, for the $k$-th instance in the batch, the model first predicts tokens for all [MASK] positions, producing an intermediate response $\mathcal{Y}^{t,k}=(\hat{y}_{i}^{t,k})_{i=1}^{L}$. We then follow the standard remasking strategy defined in existing dLLMs (e.g., selecting the top-$k$ tokens according to their confidence scores) to determine which tokens will be retained dream; dreamcoder; DiffuCoder. Subsequently and most importantly, we extract shared patterns across the generated unit test cases and leverage them to further retain additional tokens, with the detailed procedure presented in Section 3.2. After completing the procedure described above, we retain all selected tokens, while the remaining tokens at [MASK] positions are reverted back to [MASK], and the process proceeds to the next step.
@@ 3.2 Exploiting Repetitive Patterns to Accelerate UTG
We begin by analyzing the key challenges involved in leveraging such repetitive patterns, which lie in two main aspects. ❶ The first challenge is how to effectively extract shared patterns across multiple test cases. As illustrated in Figure 1, repetitive structural and syntactic patterns can be easily observed among different test cases for the same focal method. However, transforming these intuitive and abstract notions of repetitive patterns into a concrete representation-that is, determining which tokens are structurally repeated—remains a non-trivial problem. ❷ The second challenge is how to ensure sufficient diversity in the generated test cases. Achieving high test coverage relies heavily on diversity peacock2021automatic; yang2023improve, yet the process of extracting shared patterns may risk reducing it. For instance, the model might generate identical input data across multiple test cases. Designing a method that can exploit shared patterns while preserving the diversity of test data is therefore another important challenge.
@@
Extract Repetitive Patterns. We noticed Abstract Syntax Trees (ASTs) provide an objective representation of the syntactic structure of code sun2023abstract; suttichaya2022source, and each node in an AST can be accurately mapped to a specific set of tokens. Based on this property, our approach compares the ASTs of different test cases’ code and attempts to merge them as much as possible, which means identifying the nodes that are shared across different ASTs. When two or more ASTs can be merged into a non-empty tree, this indicates the existence of a repetitive pattern. We then locate the tokens in the intermediate test cases that correspond to the merged AST.
@@
Ensure Sufficient Diversity. The diversity of unit test cases largely depends on the variability of the input data constructed for the focal method, which in turn is primarily determined by literal values such as integers or floats yang2023improve. To preserve this variability, we exclude the AST nodes corresponding to such literal values from the merging process. This design ensures that even if two test cases construct similar data at an intermediate step, the corresponding tokens are not retained in a single step but are instead remasked for subsequent refinement, allowing higher diversity.
@@ 4.1 Experiment setup
Models. We evaluate our approach using two representative dLLMs, namely DiffuCoder-7B-cpGRPO DiffuCoder and Dream-v0-Instruct-7B dream. For brevity, we refer to them as DiffuCoder and Dream in the remainder of this paper.
Benchmarks. To evaluate the performance of DiffTester, we conduct experiments on the TestEval benchmark testeval. TestEval is specifically designed to assess the capability of models in unit test generation, and it comprises 210 Python programs collected from LeetCode. We denote this benchmark as TestEval-Python. To further validate the generalization ability of DiffTester across different programming languages and to avoid potential data contamination, we extend the TestEval-Python benchmark by incorporating two additional programming languages: C++ and Java. Following the construction methodology of TestEval-Python, we systematically gather corresponding C++ and Java implementations of the same set of 210 programs. We denote these benchmarks as TestEval-C++ and TestEval-Java, respectively.
@@ 4.2 Main Results
DiffTester does not compromise the maximum achievable test coverage. The highest test coverage attainable without time constraints is another important optimization target, particularly for scenarios such as vehicle control systems conrad2017testing; zhang2024visual. We find that DiffTester has little to no negative impact on the maximum test coverage and in some cases even improves it. For example, as shown in Figure 4, on the TestEval-Python benchmark with the Dream model, applying DiffTester leads to a slight increase in max test coverage.
@@ 5.1 Comparison with dLLM Acceleration for General Task
We observe that several contemporary works have also explored training-free approaches to accelerate dLLMs wei2025accelerating; ben2025accelerated; israel2025accelerating. However, these approaches typically focus on general-purpose tasks, whereas DiffTester is designed specifically for unit test generation (UTG), a domain where efficiency is particularly critical. To highlight the difference, we compare DiffTester with a representative approach, EB-Sampler ben2025accelerated. EB-Sampler accelerates dLLM inference through an Entropy-Bounded unmasking procedure, which dynamically unmasks multiple tokens under a predefined approximate error tolerance.
@@ 6.1 Unit Test Generation
Unit tests UT1; UT2 are essential for verifying the behavior of program units and serve as a foundational mechanism for early fault detection and prevention in software development. However, writing unit tests manually is highly labor-intensive, which has motivated extensive research on automated unit test generation dakhel2024effective; aniche2022effective. Traditional techniques primarily rely on search-based search_base; search-based2, random-based random_based, constraint-based constrained_based, and symbolic execution-based approaches symbolic_UT_1, with widely used tools such as EvoSuite EvoSuite, Pynguin Pynguin, and KLEE KLEE. Nevertheless, these traditional approaches often suffer from limitations such as low coverage time-consuming1.
With the rise of large language models (LLMs) achiam2023gpt, recent advances have increasingly explored their application to automated test case generation. Representative approaches include ChatTester ChatTester, TestPilot TestPilot, and ChatUnitTest ChatUnitTest. Although LLMs have significantly improved test coverage, existing LLMs used for unit test generation almost exclusively adopt a token-by-token generation paradigm zhang2025beyond, which constrains the efficiency of the unit test generation.
@@ 6.2 Diffusion Large Language Model
Diffusion language models have recently become a focal point in AI research sahoo2024simple; discreteVLA and can be categorized into continuous and discrete formulations li2025survey. Empirical evidence zhao2025d suggests that discrete diffusion language models scale more effectively to larger model sizes, which has led to the development of diffusion large language models such as LLaDA llada and Gemini Diffusion gemini_diffusion. Within this line of work, MMaDA yang2025mmada and LLaDA-V you2025llada extend dLLMs to the multimodal domain; TreeDiff zeng2025treediff explores integrating dLLMs with the structural characteristics of code; DLLM-Cache dLLM-Cache and Sparse-dLLM song2025sparse propose inference acceleration strategies based on KV-cache mechanisms; DiffuCoder DiffuCoder and LLaDA-1.5 zhu2025llada15variancereducedpreference investigate reinforcement learning tailored to dLLMs.
@@ C.1.1 Models
We primarily conducted experiments using the dLLMs DiffuCoder-7B-cpGRPO (DiffuCoder) (DiffuCoder in this paper for short) and Dream-v0-Instruct-7B dream (Dream in this paper for short).